# A parallel framework for graphical optimal transport

### v1

@@ Appendix A Proofs of Theorems
$\beta_{j}=u_{(j,k)}\odot(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\quad\text{ for }k\in N(j).$ (51)
Multiplying these relations for all $k\in N(j)$ yields
$\bigodot_{k\in N(j)}\beta_{j}=e^{\rho_{j}}\bigodot_{k\in N(j)}\left(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\right).$
Denote $v_{j}=\left(\bigodot_{k\in N(j)}(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k}))\right)^{1/|N(j)|},$ it follows that $\beta_{j}=e^{\rho_{j}/|N(j)|}v_{j},$ which together with 51 leads to the updates for $u_{(j,k)}$, $j\notin\Gamma$. Note also that the Lagrangian 49 is only bounded in $\rho_{j}$ if it holds $\beta_{j}^{\top}\mathbf{1}=1$. Thus, $e^{\rho_{j}/|N(j)|}v_{j}^{\top}{\bf 1}=1$ and the update for $\rho_{j}$ follows. ∎
Proof.
Proof of Theorem 2 Define $\Delta_{t}:=f(u^{t},\rho^{t})-f(u^{*},\rho^{*}).$ Then according to the Lemma 3, Lemma 4, there is $\Delta_{t}-\Delta_{t+1}\geq\frac{1}{22|E|}\left(\max\left\{\frac{\epsilon\Delta_{t}}{2C_{\infty}},\delta^{\prime}\right\}\right)^{2}.$ On the one hand, we have $\frac{\epsilon^{2}\Delta_{t+1}}{88|E|C_{\infty}^{2}}\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}-\left(\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\right)^{2}.$ Note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies $0\leq a_{t}\leq 1$. This implies whenever $\Delta_{t+1}\geq 0$, we have $0\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\leq 1$ for any $t>1$. Also note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies ${1}/{a_{t+1}}-{1}/{a_{t}}\geq 1$ and further implies $1{1}/{a_{t}}-{1}/{a_{1}}\geq t-$ 1. Thus,
$t\leq 1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{t}}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}.$ (52)
@@
$\Delta_{t+\tau}\leq\Delta_{t}-\frac{\delta^{\prime 2}\tau}{22|E|},\quad t,\tau\geq 0$ (53)
Fixing an number $\phi>0$, and considering using 52 for $\Delta_{t}\geq\phi$ and using 53 for $\Delta_{t}<\phi$. We minimize the sum of two estimates over the parameter $\phi\in(0,\Delta_{1}]$. Combining the above two inequalities, we get the total number of iterations satisfies
$t\leq\min_{0\leq\phi\leq\Delta_{1}}\left(1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\phi}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}+1+\frac{22|E|\phi}{\delta^{\prime 2}}\right)\leq 2+\frac{88|E|C_{\infty}}{\epsilon\delta^{\prime}}.$ (54)
Proof.
Proof of Theorem 3 With the specific choices $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$, Lemma 6 gives
$\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}\leq\delta.$ (55)
By Theorem 2, the stopping time $t$ satisfies $t\leq 2+\frac{88|E|C_{\infty}}{\delta^{\prime}\epsilon}=\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right).$ Since each iteration of Algorithm 3 requires $\mathcal{O}(|E|d^{2})$ operations, a solution of Algorithm 3 is achieved in $\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$ On the other hand, Algorithm 4 takes $\mathcal{O}(|E|d^{2})$ (see Lemma 7 in Altschuler et al. (2017)). Therefore, the bound on total complexity follows. ∎
Appendix B Proofs of Propositions
Proof.
Proof of Proposition 1 Firstly, when $j\in\Gamma$, we have $\gamma_{j}=\mu_{j}$, and
@@ Appendix B Proofs of Propositions
$=\left(\bigodot_{k\in N(j)}q^{t}_{(j,k)})\right)^{{\frac{1}{|N(j)|}}}/e^{\rho_{j}^{t}/|N(j)|}=q^{t+1}_{j}/e^{\rho_{j}^{t}/|N(j)|},$ (58)
where we use the constraint $\sum_{k\in N(j)}\log(u_{(j,k)})=\rho_{j}\mathbf{1}$ of the dual problem 32 in the third equality, and the definition of $q_{j}^{t+1}$ in the last equality. This implies that
$\rho^{t+1}_{j}=-|N(j)|\log\left(\mathbf{1}^{\top}v^{t+1}_{j}\right)=\rho^{t}_{j}-|N(j)|\log({\bf 1}^{\top}q^{t+1}_{j})=\rho^{t}_{j}-|N(j)|\log(\|q^{t+1}_{j}\|_{1}).$ (59)
@@
Proof of Proposition 3 We introduce auxiliary variables ${{\mathbf{v}}_{i}}$, $i=1,\dots,\ell$, and rewrite the constraints 45, resulting in the following reformulation of the problem
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}$ $-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle$ (62)
subject to $\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})=\log({\mathbf{v}}_{3}\otimes\mathbf{1}),$ (63)
$\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})=\log({\mathbf{v}}_{i+1}\otimes\mathbf{1}),\quad\mbox{ for }i=3,\ldots,\ell-1$ (64)
$\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$ (65)
@@
Note here that ${\mathbf{v}}_{i}$ and ${\mathbf{u}}_{i}$ have the same dimension. The Lagrange relaxed problem of 62 is
@@
$\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\boldsymbol{\mu}_{\boldsymbol{\gamma}},\log({\mathbf{u}}_{\ell+1})\rangle+\langle{\mathbf{a}}_{2},\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})-\log({\mathbf{v}}_{3}\otimes\mathbf{1})\rangle$ (66)
$\qquad\qquad+\sum_{i=3}^{\ell-1}\langle{\mathbf{a}}_{i},\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})-\log({\mathbf{v}}_{i+1}\otimes\mathbf{1})\rangle+\langle{\mathbf{a}}_{\ell},\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})-\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1})\rangle,$
@@
where ${\mathbf{a}}_{i}$ are newly introduced Lagrange variables. Maximizing with respect to ${\mathbf{v}}_{i}$ we get ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$ for $i=3,\ldots,\ell$. Maximizing with respect to ${\mathbf{v}}_{\ell+1}$, we get that $P_{\boldsymbol{\gamma}}({\mathbf{a}}_{\ell})=\boldsymbol{\mu}_{\boldsymbol{\gamma}}.$ Further, maximizing with respect to ${\mathbf{u}}_{i}$, for $i=1,\dots,\ell$, we get 48. To satisfy the constraint (63) we need ${\mathbf{u}}_{1}\odot{\mathbf{u}}_{2}={\mathbf{v}}_{3}\otimes\mathbf{1}$. Plugging in the expressions for ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$ in (48), we get the expression for ${\mathbf{a}}_{2}$ in (47) together with (46). The update of ${\mathbf{v}}_{3}$ in (47) follows from ${\mathbf{a}}_{3}=P_{c\cap c_{i}}({\mathbf{a}}_{2})$. Similarly, the remaining updates for ${\mathbf{q}}_{i}$, ${\mathbf{a}}_{i}$, and ${\mathbf{v}}_{i}$, for $i=3,\dots,\ell$, follow from the constraints ${\mathbf{v}}_{i}\odot{\mathbf{u}}_{i}={\mathbf{u}}_{\ell+1}\otimes\mathbf{1}$ and ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$. Finally, the expression for ${\mathbf{v}}_{\ell+1}$ is a consequence of $\boldsymbol{\mu}_{\gamma}=P_{\gamma}({\mathbf{a}}_{\ell})$. ∎
Appendix C Proofs of Lemmas
Proof.
@@ Appendix C Proofs of Lemmas
Proof of Lemma 1 When $j\in\Gamma\bigcap S$, the updates of $u_{(j,k)}$ would keep $\|B^{t+1}_{(j,k_{j})}\|_{1}=1$ since it matches the projection $B^{t+1}_{(j,k_{j})}\mathbf{1}$ and the true marginal $\mu_{j}$. When $j\in(V\backslash\Gamma)\bigcap S$, by proposition 1, the updates of $u_{(j,k)}$ are equivalent to 34 in Theorem 1, thus they satisfy 51, which lead to $\|B_{(j,k)}^{t+1}\|=\mathbf{1}^{\top}\beta^{t+1}_{j}.$ Since $\mathbf{1}^{\top}\beta^{t+1}_{j}=1$ holds, we get $\|B_{(j,k)}^{t+1}\|=1.$
Proof.
Proof of Lemma 2 By 33, for $j\in\Gamma$, following the logic in Dvurechensky et al. (2018, Lemma 1), there is
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}.$ (67)
By 34, for $j\in V\backslash\Gamma$, following Kroshnin et al. (2019, Lemma 4), there is
$\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}+\frac{1}{|N(j)|}\sum_{i\in N(j)}\|C_{(j,i)}\|_{\infty}.$ (68)
Combining these results gives the bound. ∎
Proof.
Proof of Lemma 3 By the gradient inequality for convex functions it holds that
$~{}~{}~{}~{}f(u^{t},\rho^{t})-f(u^{*},\rho^{*})=f(\lambda^{t},\rho^{t})-f(\lambda^{*},\rho^{*})$ (69)
@@
$=\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},\frac{1}{\epsilon}B_{(j,k_{j})}^{t}\mathbf{1}-\frac{1}{\epsilon}\mu_{j}\rangle$
$+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\langle\lambda_{(j,k)}^{t}-\lambda_{(j,k)}^{*},\frac{1}{\epsilon}q_{(j,k)}^{t}\rangle-\sum_{j\notin\Gamma}(\rho_{j}^{t}-\rho_{j}^{*})$
@@
$+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~{},~{}q_{(j,k)}^{t}\right\rangle.$
Here, we have used that $\langle\mathbf{1},q_{(j,k)}^{t}\rangle=1$, and thus $-\rho_{j}^{t}+\rho_{j}^{*}=\langle-\rho_{j}^{t}\mathbf{1}+\rho_{j}^{*}\mathbf{1},q_{(j,k)}^{t}\rangle$. We bound the first sum as in Fan et al. (2022a, proof of Lemma 2) (compare also Dvurechensky et al. (2018, proof of Lemma 2)), which leads to
$\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\rangle\leq 2C_{\infty}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}.$ (70)
To bound the second sum in the final expression of 69 note that for any $j\notin\Gamma,$
@@
$~{}~{}~{}\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~{},~{}q_{(j,k)}^{t}\right\rangle$ (71)
$=\frac{1}{\epsilon}\sum_{k\in N(j)}\left\langle\lambda_{(j,k)}^{t}-\lambda_{(j,k)}^{*}~{},~{}q_{(j,k)}^{t}-\bar{q}_{j}^{t}\right\rangle\leq\frac{2C_{\infty}}{\epsilon}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}.$ (72)
Here, the last step follows with the same approach as for 70 (see also Kroshnin et al. (2019, proof of Lemma 5)). Finally, combining 69, 70 and 72 completes the proof. ∎
Proof.
Proof of Lemma 4 Using the relations in Algorithm 3 and Lemma 1, we get
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})=\sum_{j\in\Gamma}(\log u_{(j,k_{j})}^{t+1}-\log u_{(j,k_{j})}^{t})^{\top}\mu_{j}+\sum_{j\notin\Gamma}(\rho_{j}^{t+1}-\rho_{j}^{t})$ (73)
$=$ $\sum_{j\in\Gamma}\left({{\mathcal{H}}}\left(\mu_{j}|B_{(j,k_{j})}^{t}\mathbf{1}\right)+1\right)-\sum_{j\notin\Gamma}|N(j)|\log\left(\|q_{j}^{t+1}\|_{1}\right).$ (74)

### v2

@@ 7 Proofs of Theorems
(37) $\beta_{j}=u_{(j,k)}\odot(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})),\quad\text{ for }k\in N(j).$
Multiplying for all $k\in N(j)$ yields $\bigodot_{k\in N(j)}\beta_{j}=e^{\rho_{j}}\bigodot_{k\in N(j)}\left(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k})\right)$. Denote $v_{j}=\left(\bigodot_{k\in N(j)}(K_{(j,k)}(u_{(k,j)}\odot\gamma_{k}))\right)^{1/|N(j)|},$ it follows that $\beta_{j}=e^{\rho_{j}/|N(j)|}v_{j},$ which together with Eq. 37 leads to the updates for $u_{(j,k)}$, $j\notin\Gamma$. Note also that the Lagrangian Eq. 36 is only bounded in $\rho_{j}$ if it holds $\beta_{j}^{\top}\mathbf{1}=1$. Thus, $e^{\rho_{j}/|N(j)|}v_{j}^{\top}{\bf 1}=1$ and the update for $\rho_{j}$ follows.
Proof 7.2 (Proof of Theorem 3.8).
Define $\Delta_{t}:=f(u^{t},\rho^{t})-f(u^{*},\rho^{*}).$ Then according to the Lemma 3.6, Lemma 3.7, there is $\Delta_{t}-\Delta_{t+1}\geq\frac{1}{22|E|}\left(\max\left\{\frac{\epsilon\Delta_{t}}{2C_{\infty}},\delta^{\prime}\right\}\right)^{2}.$ On the one hand, we have $\frac{\epsilon^{2}\Delta_{t+1}}{88|E|C_{\infty}^{2}}\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}-\left(\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\right)^{2}.$ Note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies $0\leq a_{t}\leq 1$. This implies whenever $\Delta_{t+1}\geq 0$, we have $0\leq\frac{\epsilon^{2}\Delta_{t}}{88|E|C_{\infty}^{2}}\leq 1$ for any $t>1$. Also note that $0\leq a_{t+1}\leq a_{t}-a_{t}^{2}$ implies ${1}/{a_{t+1}}-{1}/{a_{t}}\geq 1$ and further implies ${1}/{a_{t}}-{1}/{a_{1}}\geq t-1$. Thus,
(38) $t\leq 1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{t}}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}.$
@@
(39) $\Delta_{t+\tau}\leq\Delta_{t}-\frac{\delta^{\prime 2}\tau}{22|E|},\quad t,\tau\geq 0.$
Fixing an number $\phi>0$, and considering using Eq. 38 for $\Delta_{t}\geq\phi$ and using Eq. 39 for $\Delta_{t}<\phi$. We minimize the sum of two estimates over the parameter $\phi\in(0,\Delta_{1}]$. Combining the above two inequalities, we get the total number of iterations satisfies
(40) $t\leq\min_{0\leq\phi\leq\Delta_{1}}\left(1+\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\phi}-\frac{88|E|C_{\infty}^{2}}{\epsilon^{2}\Delta_{1}}+1+\frac{22|E|\phi}{\delta^{\prime 2}}\right)\leq 2+\frac{88|E|C_{\infty}}{\epsilon\delta^{\prime}}.$
Proof 7.3 (Proof of Theorem 3.11).
With the specific choices $\epsilon=\frac{\delta}{4|E|\log(d)}$ and $\delta^{\prime}=\frac{\delta}{8C_{\infty}}$, Lemma 3.10 gives
(41) $\sum_{(j,k)\in E}{\langle}C_{(j,k)},\widehat{B}_{(j,k)}{\rangle}-\sum_{(j,k)\in E}{\langle}C_{(j,k)},B^{*}_{(j,k)}{\rangle}\leq\delta.$
By Theorem 3.8, the stopping time $t$ satisfies $t\leq 2+\frac{88|E|C_{\infty}}{\delta^{\prime}\epsilon}=\mathcal{O}\left(\frac{|E|^{2}C_{\infty}^{2}\log(d)}{\delta^{2}}\right).$ Since each iteration of Algorithm 2 requires $\mathcal{O}(|E|d^{2})$ operations, a solution of Algorithm 2 is achieved in $\mathcal{O}\left(\frac{|E|^{3}C_{\infty}^{2}d^{2}\log(d)}{\delta^{2}}\right).$ On the other hand, Algorithm 3 takes $\mathcal{O}(|E|d^{2})$ (see Lemma 7 in [altschuler2017near]). Therefore, the bound on total complexity follows.
8 Proofs of Propositions
Proof 8.1 (Proof of Proposition 3.3).
Firstly, when $j\in\Gamma$, we have $\gamma_{j}=\mu_{j}$, and
@@ 8 Proofs of Propositions
$=\left(\bigodot_{k\in N(j)}q^{t}_{(j,k)})\right)^{{\frac{1}{|N(j)|}}}/e^{\rho_{j}^{t}/|N(j)|}=q^{t+1}_{j}/e^{\rho_{j}^{t}/|N(j)|},$
where we use the constraint $\sum_{k\in N(j)}\log(u_{(j,k)})=\rho_{j}\mathbf{1}$ of the dual problem Eq. 23 in the third equality, and the definition of $q_{j}^{t+1}$ in the last equality. This implies that
$\rho^{t+1}_{j}=-|N(j)|\log\left(\mathbf{1}^{\top}v^{t+1}_{j}\right)=\rho^{t}_{j}-|N(j)|\log({\bf 1}^{\top}q^{t+1}_{j})=\rho^{t}_{j}-|N(j)|\log(\|q^{t+1}_{j}\|_{1}).$
@@
We introduce auxiliary variables ${\mathbf{v}}_{i}$}, $i=1,\dots,\ell$, and rewrite the constraints Eq. 32, resulting in the following reformulation of the problem
(42) $\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}$ $-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\bm{\mu}_{\bm{\gamma}},\log({\mathbf{v}}_{\ell+1})\rangle$
(43) subject to $\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})=\log({\mathbf{v}}_{3}\otimes\mathbf{1}),$
(44) $\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})=\log({\mathbf{v}}_{i+1}\otimes\mathbf{1}),\quad\mbox{ for }i=3,\ldots,\ell-1$
(45) $\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})=\log({\mathbf{v}}_{\ell+1}\otimes\mathbf{1}).$
@@
Note here that ${\mathbf{v}}_{i}$ and ${\mathbf{u}}_{i}$ have the same dimension. The Lagrange relaxed problem of Eq. 42 is
@@
(46) $\max_{\begin{subarray}{c}{\mathbf{u}}_{i},i=1,\ldots,\ell\\ {\mathbf{v}}_{i},i=3,\ldots,\ell+1\end{subarray}}-\sum_{i=1}^{\ell}{\langle}{\mathbf{k}}_{i},{\mathbf{u}}_{i}{\rangle}+\langle\bm{\mu}_{\bm{\gamma}},\log({\mathbf{u}}_{\ell+1})\rangle+\langle{\mathbf{a}}_{2},\log({\mathbf{u}}_{1})+\log({\mathbf{u}}_{2})-\log({\mathbf{v}}_{3}\!\otimes\!\mathbf{1})\rangle$
$\ +\!\sum_{i=3}^{\ell-1}\langle{\mathbf{a}}_{i},\log({\mathbf{v}}_{i})+\log({\mathbf{u}}_{i})-\log({\mathbf{v}}_{i+1}\!\otimes\!\mathbf{1})\rangle+\langle{\mathbf{a}}_{\ell},\log({\mathbf{v}}_{\ell})+\log({\mathbf{u}}_{\ell})-\log({\mathbf{v}}_{\ell+1}\!\otimes\!\mathbf{1})\rangle,$
@@
where ${\mathbf{a}}_{i}$ are newly introduced Lagrange variables. Maximizing with respect to ${\mathbf{v}}_{i}$ we get ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$ for $i=3,\ldots,\ell$. Maximizing with respect to ${\mathbf{v}}_{\ell+1}$, we get that $P_{\bm{\gamma}}({\mathbf{a}}_{\ell})=\bm{\mu}_{\bm{\gamma}}.$ Further, maximizing with respect to ${\mathbf{u}}_{i}$, for $i=1,\dots,\ell$, we get Eq. 35. To satisfy the constraint (43) we need ${\mathbf{u}}_{1}\odot{\mathbf{u}}_{2}={\mathbf{v}}_{3}\otimes\mathbf{1}$. Plugging in the expressions for ${\mathbf{u}}_{1}$ and ${\mathbf{u}}_{2}$ in (35), we get the expression for ${\mathbf{a}}_{2}$ in (34) together with (33). The update of ${\mathbf{v}}_{3}$ in (34) follows from ${\mathbf{a}}_{3}=P_{c\cap c_{i}}({\mathbf{a}}_{2})$. Similarly, the remaining updates for ${\mathbf{q}}_{i}$, ${\mathbf{a}}_{i}$, and ${\mathbf{v}}_{i}$, for $i=3,\dots,\ell$, follow from the constraints ${\mathbf{v}}_{i}\odot{\mathbf{u}}_{i}={\mathbf{u}}_{\ell+1}\otimes\mathbf{1}$ and ${\mathbf{a}}_{i}=P_{c\cap c_{i}}({\mathbf{a}}_{i-1})$. Finally, the expression for ${\mathbf{v}}_{\ell+1}$ is a consequence of $\bm{\mu}_{\gamma}=P_{\gamma}({\mathbf{a}}_{\ell})$.
9 Proofs of Lemmas
Proof 9.1 (Proof of Lemma 3.4).
@@ 9 Proofs of Lemmas
When $j\in\Gamma\bigcap S$, the updates of $u_{(j,k)}$ keep $\|B^{t+1}_{(j,k_{j})}\|_{1}=1$, since it matches the projection $B^{t+1}_{(j,k_{j})}\mathbf{1}$ and the true marginal $\mu_{j}$. When $j\in(V\backslash\Gamma)\bigcap S$, by Proposition 3.3, the updates of $u_{(j,k)}$ are equivalent to Eq. 25 in Theorem 3.1, thus they satisfy Eq. 37, which lead to $\|B_{(j,k)}^{t+1}\|=\mathbf{1}^{\top}\beta^{t+1}_{j}.$ Since $\mathbf{1}^{\top}\beta^{t+1}_{j}=1$ holds, we get $\|B_{(j,k)}^{t+1}\|=1.$
Proof 9.2 (Proof of Lemma 3.5).
By Eq. 24, for $j\in\Gamma$, following the logic in [dvurechensky2018computational, Lemma 1], there is
(47) $\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}.$
By Eq. 25, for $j\in V\backslash\Gamma$, following [kroshnin2019complexity, Lemma 4], there is
(48) $\max_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}-\min_{i=1,\dots,d}(\lambda_{(j,k)}^{t})_{i}\leq\|C_{(j,k)}\|_{\infty}+\frac{1}{|N(j)|}\sum_{i\in N(j)}\|C_{(j,i)}\|_{\infty}.$
Combining these results gives the bound.
Proof 9.3 (Proof of Lemma 3.6).
By the gradient inequality for convex functions it holds that
$f(u^{t},\rho^{t})-f(u^{*},\rho^{*})=f(\lambda^{t},\rho^{t})-f(\lambda^{*},\rho^{*})$
@@
$=\!\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}\!-\!\lambda_{(j,k_{j})}^{*},\frac{1}{\epsilon}B_{(j,k_{j})}^{t}\mathbf{1}\!-\!\frac{1}{\epsilon}\mu_{j}\rangle\!+\!\sum_{j\notin\Gamma}\sum_{k\in N(j)}\!\!\!\langle\lambda_{(j,k)}^{t}\!-\!\lambda_{(j,k)}^{*},\frac{1}{\epsilon}q_{(j,k)}^{t}\rangle-\!\sum_{j\notin\Gamma}(\rho_{j}^{t}\!-\!\rho_{j}^{*})$
@@
$\quad+\sum_{j\notin\Gamma}\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~,~q_{(j,k)}^{t}\right\rangle.$
Here, we have used that $\langle\mathbf{1},q_{(j,k)}^{t}\rangle=1$, and thus $\rho_{j}^{t}+\rho_{j}^{*}=\langle-\rho_{j}^{t}\mathbf{1}+\rho_{j}^{*}\mathbf{1},q_{(j,k)}^{t}\rangle$. We bound the first sum as in [fan2021complexity, proof of Lemma 2] (compare also [dvurechensky2018computational, proof of Lemma 2]), which leads to
(49) $\sum_{j\in\Gamma}\langle\lambda_{(j,k_{j})}^{t}-\lambda_{(j,k_{j})}^{*},B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\rangle\leq 2C_{\infty}\sum_{j\in\Gamma}\|B_{(j,k_{j})}^{t}\mathbf{1}-\mu_{j}\|_{1}.$
To bound the second sum in the final expression of 9.3 note that for any $j\notin\Gamma,$
@@
(50) $~~~\sum_{k\in N(j)}\left\langle\frac{1}{\epsilon}\lambda_{(j,k)}^{t}-\frac{1}{|N(j)|}\rho_{j}^{t}\mathbf{1}-\frac{1}{\epsilon}\lambda_{(j,k)}^{*}-\frac{1}{|N(j)|}\rho_{j}^{*}\mathbf{1}~,~q_{(j,k)}^{t}\right\rangle$
(51) $=\frac{1}{\epsilon}\sum_{k\in N(j)}\left\langle\lambda_{(j,k)}^{t}-\lambda_{(j,k)}^{*}~,~q_{(j,k)}^{t}-\bar{q}_{j}^{t}\right\rangle\leq\frac{2C_{\infty}}{\epsilon}\sum_{k\in N(j)}\|B_{(j,k)}^{t}\mathbf{1}-\bar{q}_{j}^{t}\|_{1}.$
Here, the last step follows with the same approach as for Eq. 49 (see also [kroshnin2019complexity, proof of Lemma 5]). Finally, combining 9.3, Eq. 49 and Eq. 51 completes the proof.
Proof 9.4 (Proof of Lemma 3.7).
Using the relations in Algorithm 2 and Lemma 3.4, we get
$f(u^{t},\rho^{t})-f(u^{t+1},\rho^{t+1})=\sum_{j\in\Gamma}(\log u_{(j,k_{j})}^{t+1}-\log u_{(j,k_{j})}^{t})^{\top}\mu_{j}+\sum_{j\notin\Gamma}(\rho_{j}^{t+1}-\rho_{j}^{t})$
$=$ $\sum_{j\in\Gamma}\left({{\mathcal{H}}}\left(\mu_{j}|B_{(j,k_{j})}^{t}\mathbf{1}\right)+1\right)-\sum_{j\notin\Gamma}|N(j)|\log\left(\|q_{j}^{t+1}\|_{1}\right).$